Suppose a​ five-year,$1,000 bond with annual coupons has a price of $901.51 and a yield to maturity of 6.4%.What is the​ bond's coupon​ rate?
Answer #1
Bond's price is PV of redemption price and PV of coupon payments
Bond Price 901.51
Redemption value 1000
Time 5
coupon rate R
YTM 6.40%
PV of redemption price =1000/(1+6.4%)^5 733.3172
So rest PV belongs to Coupon payments =901.51-733.317
168.193
PV of annuity for making pthly payment
P = PMT x (((1-(1 + r) ^- n)) / i)
Where:
P = the present value of an annuity stream
PMT = the dollar amount of each annuity payment
r = the effective interest rate (also known as the discount rate)
i=nominal Interest rate
n = the number of periods in which payments will be made
PV of coupon payments =PMT x (((1-(1 + r) ^- n)) / i)
168.193 =PMT x (((1-(1 + 6.4%) ^- 5)) / 6.4%)
168.193 =PMT x 4.1669
Annual Coupon rate= 168.193/4.1669
Annual Coupon amount= 40.36406
Annual Coupon RATE= 40.364/1000
Annual Coupon RATE= 4.04%